Operations Report – February 25th

Crew 189 Operations Report 25Feb2018

SOL: 7

Name of person filing report: Gabriel PAYEN

Non-nominal systems:

-Pipe leak at the low desk

Notes on non-nominal systems:

-It seems that the pipe leak comes from the kitchen faucet.

Generator (hours run): Turned on at 9:AM and turned off at 6:00 PM

Solar— SOC % (Before generator is run at night) : 96%

Diesel – 55 %

Propane – 32 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –

Water (trailer) – –

Water (static) – –

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 131 439.1

Toilet tank emptied: NO

ATV’s Used: None

Oil Added –

ATV Fuel Used: Gals

Hours the ATVs were Used today: –

Notes on ATVs:

Deimos rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where –

General notes and comments: –

Summary of internet: –

Summary of suits and radios: –

Summary of Hab operations: –

Summary of GreenHab operations: –

Summary of ScienceDome operations: –

Summary of RAMM operations: –

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support: –

Sol Summary – February 25th

Crew 189 Sol Summary Report 25 Feb 2018

Sol 7

Summary Title: Rest Day

Author: Louis Mangin

Mission Status: All nominal

Sol Activity Summary:
We decided to declare two rest days this year because all the lack of
sleep we accumulated last year. No fixed hour of wake up for today,
but everybody was busy by 11:00 a.m. This day was supposed to be free
and dedicated to long run tasks or personal time. In the afternoon, we
split in different team, working in the greenhab to rearrange it,
cooking, producing an encouragement video for school’s student
selected for a TV show. It was also the time for a first individual
debriefing about the sport sessions I organize.

Look Ahead Plan: Change of LOAC’s battery, exploration of White Canyon
in EVA. Experiments repair, backpack change of battery, emergency
procedures inside training.

Anomalies in work: Water leak in the Hab Lower deck

Weather: Mostly sunny, cold.

Crew Physical Status: Fine.

EVA: –

Reports to be file:
Commander
Journalist
Engineering
Greenhab Officer

Support Requested: –

Commander Report – February 25th

Crew 189 Commander Report 25Feb2018

Sol 7

Title: First day off

Dear Earth,

Last year misadventures made Louis and I declare one day off every week.

Sol7: we have been there for an entire week now, and it was our first day off (really appreciated by everyone). Since our arrival, we had been doing our thirty-minutes physical training every single Sol, at 7 AM before breakfast, and almost an EVA every morning.

With Louis, we decided this year to make one day off per week in order to have some rest, to have fun altogether and then to work better the other days. This morning was supposed to be without any early alarms and each crewmember could sleep as they wanted. Nevertheless, I had quite a noisy and (really) boring alarm at 5am. My dear greenhab officer, who is sleeping in the room next to mine, snored all night long and this horrible noise woke me up. As I could not manage to sleep again, I had breakfast in this quiet station (I have to say living with six men is sometimes really noisy). I spent a really nice moment, looking behind the windows to see the Martian sunset. As Gabriel failed cooking the bread yesterday evening (he used gluten instead of flour in his bread), he decided to cook another one just after and I was the first to taste it: it was an amazing breakfast! Not really awake, I enjoyed this moment and decided to have some rest after this really short night.

I spent two hours in my bedroom, enjoying the blue sky and the sunny light from my window. When the other crewmembers woke up, we talked and spent time chilling altogether. Jérémy and Benoit went to the greenhab in order to re-organize it. Gabriel had his first shower (2,3 gallons) and I filmed again a pep talk video for our fellow colleagues on Earth who will take part into a famous French TV game.

Louis worked on the video-editing all day (hopefully he knows how to do it and enjoys this).

Finally, our Sundays on Mars are made of late wake up, no early workout, kind of brunch and resting activities (reading, watching movies, playing games, cooking…). We are all together on Mars for one week already and I think the crew stays in a good mood and is feeling well on this new planet. (Let’s pray my fellow neighbor will not snore tonight…)

Ad Astra!

Victoria Da-Poian

Commander of the Sunday-chilling Crew 189

Cheers

Operations Report – February 21st

Crew 189 Operations Report 21/02/2018
SOL: 3

Name of person filing report: Gabriel PAYEN

Non-nominal systems: Pipe leak, Fans on backpack 2 & 6, roof heater, light bulb in the bathroom, hesitation about burnable/unburnable trashes.

-Pipe :
The pipe in the low desk still leaks

-Roof Heater :
It’s still difficult to control the heater

-Backpacks :
All engineering checks were nominals, but during the EVA fans of backpack 2 and 6 turned out of service and both helmets were very foggy.
Backpacks 2 and 6 joined backpacks 3 and 4 to the list of backpacks the crew need to fix. The crew was supposed to fix them today but the we were all very busy. I hope we’ll begin tomorrow.

-Light bulb :
At 5:00 PM, a crew member was unable to light on the bulb in the bathroom. It seems that it is broken.

-Trashes :
During the EVA engineering check, the EVA members had to but five wastes bags into the outside trashes but “Burnable” and “Unburnable” weren’t indicated. Therefore, they randomly put the wastes into the trashes, which is quite pity considering we do a selective sorting for wastes.

Generator (hours run): Turned of at 12:20 and turned on at 18:08

Solar— SOC % (Before generator is run at night) –

Diesel – 70 %

Propane – 39 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 7 gallons.

Water (trailer) – 0 gallons.

Water (static) – 300 gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 131 293.2

Toilet tank emptied: YES

ATV’s Used: 350.1 350.2

Oil Added : No

ATV Fuel Used: Gals

Hours the ATVs were Used today: 3 hours

Notes on ATVs: Because of the cold night, it took a few minutes to turn them on.

Deimos rover used: NO
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: NO
Hours:
Beginning charge:
Ending charge:
Currently charging:

Opportunity rover used: YES
Hours: 16.8
Beginning charge: 100%
Ending charge: 75%
Currently charging: YES

Curiosity rover used: NO
Hours:
Beginning charge:
Ending charge:
Currently charging:

HabCar used and why, where : NO

General notes and comments: –

Summary of internet: CapCom Bernard wanted some information about the Hab Router. I gave him it by email.

Summary of suits and radios: Backpack 2 & 6 put aside.

Summary of Hab operations: –

Summary of GreenHab operations: Jeremy (GreenHab officer) began to set the new shade with the help of other crew members.

Summary of ScienceDome operations: –

Summary of RAMM operations: –

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support:

Bulb :
Is it possible to have a new light bulb for the bathroom ?
Brand, model and specifications of the old one : Lights of America – 2814 S – 120V 60Hz 14W 255mA

Pipe leak :
Does a map of all Hab pipes exist ? I still don’t find were the leak comes from. I thought it could be a cooling or heating water, therefore I checked both heaters (roof and wall) today, and they are OK.
The crew is out of ideas considering this leak (I also checked taps yesterday), therefore such a map can help me.

Trashes :
The crew needs indications about trashes to know exactly in which ones burnable and unburnable wastes have to be put into.

Operations Report – February 21st

Crew 189 Operations Report 21Feb2018
SOL: 2

Name of person filing report: PAYEN Gabriel

Non-nominal systems: Backpack 3

WiFi

Solar panels < 70% without generator running

Notes on non-nominal systems:

Backpack 3 :

All engineering checks were nominals, but during the EVA both fans turned out of service. Also, the necklace which holds the helmet was not stretched enough so that the helmet fell on the neck of Gabriel during half of the EVA.

Backpack 3 joined backpack 4 to the list of backpacks the crew need to fix. We’ll certainly begin to fix them tomorrow.

WiFi :

Only 15 Mo were left this morning (immediately after the very connection of the day were established). Yet, the leak is unknow. I reported the problem this morning and the crew have communicated with CapCom to find a solution.

Solar panels :

I had to go out of the SIM to turn on the generator at 6:20 PM. Solar SOC was at 53%, I hope batteries are not injured.

Generator (hours run): Turned of at 12:00 and turned on at 6:20PM.

Solar— SOC % (Before generator is run at night) /!\ 53%

Diesel – 70%

Propane – 40 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 7 gallons.

Water (trailer) – 0 gallons.

Water (static) – 65% gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 0.26m

Toilet tank emptied: NO

ATV’s Used: NO

Deimos rover used: YES

Hours: 109.1

Beginning charge: 100%

Ending charge: 100% (the EVA team doesn’t travel a lot)

Currently charging: YES

Spirit rover used: NO

Hours: 20.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: NO

Opportunity rover used: YES

Hours: 16.6

Beginning charge: 100%

Ending charge: 45%

Currently charging: YES

Curiosity rover used: NO

Hours: 100%

Beginning charge: –

Ending charge: –

Currently charging: NO

General notes and comments: –

Summary of internet:

-Still seeking the leak.

Summary of suits and radios:

-Backpack 3 is put aside. Starting fixing backpack 3 and 4 tomorrow.

Summary of Hab operations:

-Two thermometers and the shape for the GreenHab received in the airlock. Jeremy (GreenHab officer) will set it up tomorrow.

-The pipe in the low desk seems not to flow anymore. All water devices in the Hab and GreenHab were tested one by one and none trigger the leak.

-Considering a Dr Rupert’s request, a mattress will be put in the airlock between 8PM and 9 PM.

Summary of GreenHab operations: –

Summary of ScienceDome operations: –

Summary of RAMM operations: –

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support: –

GreenHab Report – February 21st

GreenHab Report

Jérémy Auclair

21 February 2018

Environmental control:

Ambient with door closed

Fan during 30 minutes

Shade cloth on

Average temperatures: I put the new thermometers in the greenhab today and the temperature was about 16 degrees Celsius at 5pm, went up to 30 °C at 8pm after we worked in the GreenHab

Hours of supplemental light: from 6pm to 11pm

Changes to crops: one cucumber that was not ripes fell down while changing the shade

Daily water usage for crops: 13 gallons

Time(s) of watering for crops: 10:30apm and 7:30pm

Aquaponics: None

Narrative: Today, the greenhab had many visitors. We installed the new shades (80% shading) and it was a hassle, it took us a lot of time and was very complicated! Almost all the crew participated in this installation. The commanders and the astronomer helped a lot and without them I would not have been able to do it! Thanks to this installation I noticed that many tomatoes plants were bending and with Benoit we used many wooden sticks. We changed the organisation of the GreenHab quite a lot. I still have many work to do to rearrange all this but, at least, the new cover is partially installed!

Support/supplies needed: Wooden sticks (again please)

EVA Report

Hello again CapCom,

Here is today’s EVA report:

EVA Report:
EVA #3 Crew members: Louis Mangin (EVA leader), Benoit Floquet, Laurent Bizien, Alexandre Martin

Author: Louis Mangin
Location: Around the Hab 12S-518100E-4250700N
Near Pooh’s Corner: 12S-518900E-4250900N
Cactus road: 12S-52200E-4252700N

Time: departure at 9:30 p.m. Duration: 2 hours

Purposes:
– Deploy the LOAC after reparation
– Check MegaARES status
– Explore around Cactus Road
– Shoot Media photos

Narrative:
We started by getting a lot of gear out of the engineering airlock: wastes, LOAC and tools. We then struggled to start the ATVs because of the coldness. When arrived on site, I had fog on my helmet because of the walk carrying the device, so that I didn’t see exactly chat happened, but while plugging the power wire, the power system overheated, and wires melted. I then disconnected the battery, but it was already too late. We evaluated then that only the power system was touched and apparently were right to guess so. Anyway, we packed everything back and went back to the hab to bring the sensor back. At the same time Benoit had checked the MegaARES and everything was fine there. After letting the broken LOAC in the engineering airlock, we headed to Cactus Road. We then climbed a hill, went to an entrance of Candor Chasma and shoot some group photos with our school’s banner. The fog intensified for the ones wearing the first version of the suits, and it was pretty cold, so that we headed back to the Hab early.

9:30 a.m.: EVA#3 left Airlock
9:40 a.m.: EVA left the Hab using two ATVs and one rover
9:45 a.m.: Arrival at place of interest
10:00 a.m.: Failure due to LOAC’s power system overheating
10:20 a.m.: Back to Hab to put back the LOAC
10:30 a.m.: Leaving the Hab to Cactus Road
10:40 a.m.: Exploration of Cactus Road surroundings
11:15 a.m.: Leaving Cactus Road
11:30 a.m.: EVA End

Louis Mangin

Sol Summary – February 21st

Crew 189 Sol Summary Report 21 Feb 2018

Sol 3

Summary Title: First results

Author: Louis Mangin

Mission Status: All nominal

Sol Activity Summary:
Normal start: everybody is now accustomed to the physical training even if some legs are already heavy. The third EVA was supposed to deploy the LOAC but it didn’t go that well. The power system had an issue and wires melted once I pluged it in. We then brought the equipement back to the hab, and went on for the EVA, going East to do a bit of exploration near Cactus Road, and shoot photos for the school.
This afternoon was dense: after lunch break, we answered to a French journalist interview, played the first session of Gabriel’s human factor experiment, installed a new shade on the GreenHab, worked on experiments, so that we didn’t have enough time to fix the backpack’s fans. Tomorrow EVA will then be done using the new version, before we take time to take care of these (this time, I really want to do it tomorrow).

Look Ahead Plan: Third deployment try of LOAC, data gathering of MegaARES, exploration of Candor Chasma, backpack’s fans inspection, end of shade replacement in greenhab.

Anomalies in work: Broken light in the bathroom (no replacement one found). Water leak in the Hab Lower deck, broken fans in spacesuits.

Weather: Cold and Cloudy

Crew Physical Status: Fine, already a bit tired, but everybody naped.

EVA: Second

Reports to be filed:
Commander
Journalist
HSO
Engineering
Greenhab Officer
EVA

Support Requested: New light for bathroom (see Engineering report)